# A Quantitative Comparison of Pre-Trained Model Registries to Traditional Software Package Registries

### v1

@@ 5.5. $C_{3}$: Docs Quality Impacts Model Selection
5.5.1. Method
Metric 4: Documentation Quality: Prior work has examined documentation quality in many ways (Gong et al., 2023b; Tang and Nadi, 2023). We used those ideas to develop our measure of quality. The primary documentation for PTMs is called a “model card”, which is similar to the README of a GitHub repository or the landing page of an NPM or PyPI package. We considered two factors: (1) the completeness of the model card; and (2) the availability of metadata.
To measure completeness, we identified five typical sections found in highly popular PTMs such as Google’s Bet base uncased: Model Description, Limitations, How to Use, Training, and Evaluation. We scored model cards on an integer scale from 0 to 5 — to receive a score of 5, a PTM’s card needed all of these sections. To assess whether each section was present, we queried OpenAI’s ChatGPT-4 with the prompt shown in Listing LABEL:lst:prompt4docquality.
@@ 5. Evaluation: Reports on the models performance evaluation
To measure the availability of metadata, we referenced the PeaTMOSS dataset, which extracted over 20 distinct pieces of metadata if they were present in a PTM’s model card and associated configuration files. We scored PTMs on an integer scale from 0 to the maximum, i.e., the number of distinct pieces of metadata considered in the PeaTMOSS database schema. A PTM scoring perfectly in this category would possess all available metadata according to PeatMOSS.
@@
To test our interpretation of the metric, we selected the top 1000 and bottom 1000 models from PeaTMOSS’s set of 15,000 PTMs. We evaluated their documentation quality, summing an overall documentation score as a (0,1) metric that normalized and then weighted the two components equally. Then we compared the distributions using a box-and-whisker plot and statistical tests.
5.5.2. Results
@@ 6. Threats to Validity
Construct Threats are potential limitations of how we operationalized concepts. In the systematic literature review (RQ1), we manually extracted claims from papers, which might introduce potential bias to our results. As a mitigation, to improve objectivity two authors worked together on the process and the filtering of claims. In the validation of non-quantified claims, we proposed metrics that seemed suitable based on our judgment. As a mitigation, where possible we used multiple measures and leveraged previously-defined metrics.
Internal threats are those that affect cause-effect relationships. We emphasize that our approach for RQ2 is of the form: “If claim $C$ is true, then measurement $M$ should show us that…”. In each case the measurement produced the expected result. However, this result is correlative, not causative — there may be a latent variable in each case, or the causative relationship may be reversed. For example, in Figure 8 we found that better documentation correlates with greater popularity. It may be that the latent variable is performance, such that models become popular because they have good performance, and they accrue documentation because they are popular. When qualitative and quantitative claims agree, as is the case in this study, we learn both “Why?” and “How much?”.
@@ 7. Future Work
In Table 4, we identified one claim that we could not operationalize for measurement: $C_{4}$, that different deep learning-specific attributes would affect PTM selection and reuse. Although the PeaTMOSS dataset does include some of these measures in a structured format, the claim is sufficiently broad that quantifying it was beyond the scope of this study. Future work could explore the multi-variable relationship implied by this claim.
Our study provides the first systematic literature review of current knowledge about PTM reuse. Another opportunity is a systematic comparison of the software package registries associated with traditional software packages, and the software package registries associated with PTMs. Jiang et al. observed similarities and differences in the reuse processes (Jiang et al., 2023b) — how and to what extent can we measure the differences? As discussed in section 5.4.1 with respect to an analogue for PTM descendants, finding the limits of comparison (e.g., appropriate measurements) between traditional vs. PTM software package registries is an open challenge.
The rapid development of PTM technologies presents an ongoing challenge for empirical research on PTM reuse. The Hugging Face platform continues to grow exponentially (Castaño et al., 2023), the state of the art performance of models at all sizes continues to advance (pap, [n. d.]), and the tooling available to adapt and deploy these models continues to improve (Jajal et al., 2023). While datasets and tooling such as HFCommunity and PeaTMOSS support studies like ours, they also present some limitations. Innovation is needed to help empirical software engineering researchers keep up with the scale of activity and volume of data that we are seeing in the context of PTM reuse. Given Hugging Face’s dynamic growth, even data that is a few months old may not accurately reflect the current state of the platform, suggesting a need for regular snapshotting, which imposes significant storage requirements (beyond the already-substantial requirements of $>$50 TB). This highlights the need for tools that can provide real-time, incrementally-updated data to keep pace with rapid changes, ensuring that analyses remain relevant and reflective of the present situation in PTM reuse.
@@ 8. Conclusion
Pre-trained models are the motive force of the new generation of software engineering. Understanding engineers’ PTM reuse practices is crucial to optimizing and securing the process. Our systematic literature review and quantification of claims has illuminated significant aspects of PTM reuse. We also shed light on unique dynamics within the PTM reuse landscape as compared to traditional software package registries. Specifically, we observed a shorter lifespan of packages in PTM registries compared to traditional SPRs. Our findings underscore the need for research infrastructure and novel tools to support PTM reuse, adapting to the much higher turnover of popular PTM packages. We must ensure that PTM registries can meet the evolving demands of the software engineering community.
@@ 9. Data Availability
An anonymous artifact containing the results of the systematic literature review (RQ1) as well as our software and results for quantification of claims (RQ2) is available at https://github.com/anonsub1234/ptm-quantify-esem-2024.
10. Research Ethics
No human subjects were involved in the conduct of this project. We are aware of no other ethical concerns.
@@ Acknowledgments
We acknowledge support from NSF awards OAC-2107020 and OAC-2107230.

### v2

@@ 5.5. $C_{3}$: Docs Quality Impacts Model Selection
5.5.1. Methodology.
Metric 4: Documentation Quality: Prior work has examined documentation quality in many ways (Gong et al., 2023; Tang and Nadi, 2023). We used those ideas to develop our measure of quality. The primary documentation for PTMs is called a “model card”, which is similar to the README of a GitHub repository or the landing page of an NPM or PyPI package. We considered two factors: (1) the completeness of the model card; and (2) the availability of metadata.
To measure completeness, we identified five typical sections found in highly popular PTMs such as Google’s Bet base uncased: Model Description, Limitations, How to Use, Training, and Evaluation. We scored model cards on an integer scale from 0 to 5 — to receive a score of 5, a PTM’s card needed all of these sections. To check if each section was present, we queried OpenAI’s ChatGPT-4 (Listing LABEL:lst:prompt4docquality).
@@ 5. Evaluation: Reports on the models performance evaluation
To measure the availability of metadata, we referenced the PeaTMOSS dataset, which extracted over 20 distinct pieces of metadata if present in a PTM’s model card and associated configuration files. We scored PTMs on an integer scale from 0 to the maximum, i.e., the number of distinct pieces of metadata considered in the PeaTMOSS database schema. A perfect scoring PTM in this category would possess all available metadata according to PeatMOSS.
@@
To test our interpretation of the metric, we selected the top 1000 and bottom 1000 models from PeaTMOSS’s set of 15,000 PTMs. We evaluated their documentation quality, summing an overall documentation score as a (0,1) metric by normalizing the two components and weighting them equally. Then we compared the distributions using a box-and-whisker plot and statistical tests.
5.5.2. Results.
@@ 6. Threats to Validity
Construct Threats are potential limitations of how we operationalized concepts. In the systematic literature review (RQ1), we manually extracted claims from papers, which might introduce potential bias to our results. As a mitigation, to improve objectivity two authors worked together on the process and the filtering of claims. In the validation of non-quantified claims, we proposed metrics that seemed suitable based on our judgment. To mitigate, we used multiple measures and leveraged existing metrics.
Internal threats are those that affect cause-effect relationships. We emphasize that our approach for RQ2 is of the form: “If claim $C$ is true, then measurement $M$ should show us that…”. In each case the measurement produced the expected result. However, this result is correlative, not causative — there may be a latent variable in each case, or the causative relationship may be reversed.
For example, in Figure 8 we found that better documentation correlates with greater popularity. It may be that the latent variable is performance, such that models become popular because they have good performance, and they accrue documentation because they are popular. When qualitative and quantitative claims agree, as is the case in this study, we learn both “Why?” and “How much?”.
@@ 7. Future Work
Our study provides the first systematic literature review of current knowledge about PTM reuse. Another opportunity is a systematic comparison of the software package registries associated with traditional software packages, and the software package registries associated with PTMs. Jiang et al. observed similarities and differences in the reuse processes (Jiang et al., 2023a) — how and to what extent can we measure the differences? As discussed in section 5.4.1 with respect to an analogue for PTM descendants, finding the limits of comparison (e.g., appropriate measurements) between traditional vs. PTM software package registries is an open challenge.
In section 5.4.2, we identified several potential improvements for PTM platforms, particularly in the areas of version control, model lineage tracking, and recommendation systems. However, the specific attributes that these platforms should measure and analyze remain underexplored. Our findings suggest that popularity is a factor in model selection and reuse, but not the sole determinant (cf. claim $C_{4}$). Therefore, we recommend further research to uncover and quantify the deep learning-specific attributes that influence a model’s reusability and adaptability. Such research would provide valuable insights for developing more sophisticated tools and metrics for PTM platforms, enabling users to make more informed decisions when selecting models for reuse or further development.
The rapid development of PTM technologies presents an ongoing challenge for empirical research on PTM reuse. The Hugging Face platform continues to grow exponentially (Castaño et al., 2024), the state of the art performance of models at all sizes continues to advance (pap, [n. d.]), and the tooling available to adapt and deploy these models continues to improve (Jajal et al., 2024). Empirical software engineering researchers need to stay abreast of the activity and volume of data in the context of PTM reuse. Given Hugging Face’s dynamic growth, even data that is a few months old may not accurately reflect the current state of the platform, suggesting a need for regular snapshotting, which imposes significant storage requirements (beyond the already-substantial requirements of $>$50 TB). This highlights the need for tools that can provide real-time, incrementally-updated data to keep pace with rapid changes, ensuring that analyses remain relevant and reflective of the present situation in PTM reuse.
@@ 8. Conclusion
Pre-Trained Models are driving the next generation of software engineering. We must understand (and optimize) engineers’ PTM reuse practices. This systematic literature review has synthesized existing knowledge about PTM reuse in the Hugging Face registry. We also quantified certain claims, revealing unique dynamics within the PTM landscape compared to traditional software package registries. These findings motivate further research to understand and optimize the distinctive aspects of PTM registries for PTM reuse, to meet the evolving needs of the software engineering community.
@@ 9. Data Availability
An artifact is available (git, 2024), containing results of the SLR (RQ1) and code and results for claim quantification (RQ2). Source: https://github.com/PurdueDualityLab/ptm-quantify-esem-2024/.
@@ Acknowledgments
Work was supported by NSF awards #2107020 and #2107230.